OpenStudy (anonymous):

Byron’s income last year was $61,700. He made $48,000 at his plumbing job and had income from an apartment that he rented out. If Byron earned the same amount of rent each month for the entire year, approximately how much per month did he collect? a. $940 b. $1,140 c. $4,000 d. $5,140

OpenStudy (shinalcantara):

if the total income is 61,700 and 48,000 is from his plumbing job, then from the apartment he rented out would be: 61,700 - 48,000 = 13,700 since you wanted to find out the income monthly, divide it by 12 since there are 12 months in a year.. 13,700/12 = ? how much would that be?

OpenStudy (uri):

61700-48000=13700 13700/12=1141.6~ approximately